OpenStudy (anonymous):

Prove the following trigonometric identity: 2/ (√3 cos x + sin x) = sec(π/6 - x)

OpenStudy (anonymous):

\[\dfrac{2}{\sqrt3 cos x+sin x}=sec(\pi/6 -x)\] right?

OpenStudy (anonymous):

Yes. I have no idea how to do that on this site, but yes

OpenStudy (anonymous):

let go from the right hand side \[sec (\pi/6-x)=\dfrac{1}{cos(\pi/6-x)}\]

OpenStudy (anonymous):

Now, manipulate a little bit the denominator cos (pi/6 -x) = cos (pi/6) cos x+ sin(pi/6) sinx cos pi/6= sqrt3/2 sin pi/6 = 1/2 just plug on them in , simplify and you get the left hand side

OpenStudy (anonymous):

*all not on

OpenStudy (anonymous):

That looks about right. Thank you!